OpenStudy (anonymous):

Mr. Bird’s son is on a coed soccer team with Mrs. Twiddy’s daughters for a total of 36 team members. The number of girls is 3 less than double the number of boys. How many of the players are boys and how many are girls? Write a system of equations to model the problem. Solve the system by substitution, show your work.

jigglypuff314 (jigglypuff314):

Hello @Ultimate_gamer and Welcome to OpenStudy!! :) Let's first take the information that we know and make a system of equations: "total of 36" which means boys plus girls equals that much -> x + y = 36 "number of girls is 3 less than double the number of boys" -> x = 2y - 3 where x is girls and y is boys Follow so far? :) we have the system of equations of: x + y = 36 and x = 2y - 3

jigglypuff314 (jigglypuff314):

From here we can use the Substitution method: plug in x = (2y - 3) into the first equation of x + y = 36 you will get: (2y - 3) + y = 36 from there you can solve for y ! (boys) --- once you get that, you can plug in the y you got into the x = 2(y) - 3 equation and get x ! (girls) that will give you the two answers you need :)